# There Was Never a Bottleneck in Concept Bottleneck Models

### v1

@@ Abstract
Deep learning representations are often difficult to interpret, which can hinder their deployment in sensitive applications. Concept Bottleneck Models (CBMs) have emerged as a promising approach to mitigate this issue by learning representations that support target task performance while ensuring that each component predicts a concrete concept from a predefined set. In this work, we argue that CBMs do not impose a true bottleneck: the fact that a component can predict a concept does not guarantee that it encodes only information about that concept. This shortcoming raises concerns regarding interpretability and the validity of intervention procedures. To overcome this limitation, we propose Minimal Concept Bottleneck Models (MCBMs), which incorporate an Information Bottleneck (IB) objective to constrain each representation component to retain only the information relevant to its corresponding concept. This IB is implemented via a variational regularization term added to the training loss. As a result, MCBMs support concept-level interventions with theoretical guarantees, remain consistent with Bayesian principles, and offer greater flexibility in key design choices.
@@ 1 Introduction
Deep learning (DL) models have become the de facto standard for solving a wide range of tasks, primarily due to their ability to uncover complex patterns in data. However, these patterns often go beyond human comprehension, leading to DL models being perceived as black boxes—their predictions may be accurate, but they are typically difficult to interpret. Despite this, interpretability remains a crucial requirement in several domains, including but not limited to healthcare [1, 2, 3], finance [4, 5], and autonomous driving [6, 7].
Most DL models operate by learning data representations—compressed versions of the input that retain essential information for tackling specific tasks of interest [8]. Concept Bottleneck Models (CBMs) aim to define these representations based on a set human-understandable concepts [9]. Given a task $\bm{y}$ and a set of concepts $\bm{c}=\{c_{j}\}_{j=1}^{m}$, CBMs obtain a representation $\bm{z}$ and are trained to jointly optimize the two following objectives:
  1. (i)
    The task $\bm{y}$ must be perfectly solved.
  2. (ii)
    Each concept $c_{j}$ must be predictable by using only a subset $z_{j}\in\bm{z}$.
As a result of the second objective, CBMs are purported to offer: (i) increased interpretability of the representation space, and (ii) the ability to intervene a concrete concept $c_{j}$ by modifying $z_{j}$ and propagating these changes to the predictions.
@@
However, a phenomenon called information leakage has been observed in CBMs [10, 11], whereby $\bm{z}$ encodes information that is not attributable to the defined concepts. This issue poses two major concerns: (i) it undermines the interpretability of the learned representations, and (ii) it compromises the validity of interventions—modifying $z_{j}$ may alter not only the associated concept $c_{j}$, but also other unintended factors encoded in $z_{j}$. We argue that this issue stems from a fundamental limitation in the current formulation of CBMs: the absence of an explicit Information Bottleneck (IB) [12] that actively constrains $z_{j}$ to exclude information unrelated to $c_{j}$. While the second objective in CBMs encourages each $z_{j}$ to retain $c_{j}$ in its entirety, it does not enforce that $z_{j}$ captures only information about $c_{j}$. In the worst-case scenario, $z_{j}$ could encode the entire input $\bm{x}$ and still satisfy this objective.
We propose a solution to the aforementioned problem by introducing an IB in $z_{j}$ in such a way that it does not only preserve all the information of $c_{j}$, but it also contains information only about $c_{j}$. We call these models Minimal Concept Bottleneck Models (MCBMs), since $z_{j}$ is trained to be a minimal sufficient statistic of $c_{j}$ [13, 14]. This contrasts with traditional CBMs, where $z_{j}$ is optimized to be merely a sufficient statistic of $c_{j}$. In other words, while CBMs simply require that $z_{j}$ preserves all information about $c_{j}$, MCBMs require that $z_{j}$ preserves only all the information about $c_{j}$. We implement the IB constraint via an additional loss term derived using variational approximations. This formulation enables faithful concept interventions without violating Bayesian principles and offers greater flexibility in key choices of the model.
2 Related Work
Information Leakage in Concept Bottleneck Models
This phenomenon has been observed in CBMs trained either sequentially or jointly [10, 11], and it arises when the learned representation $\bm{z}$ encodes information that is not present in the concept set $\bm{c}$. Such leakage negatively impacts both the interpretability and the intervenability of the model. As shown in [15], this issue arises when the Markovian assumption fails—i.e., when the target $\bm{y}$ is not fully determined by the concept set $\bm{c}$, or equivalently, $p(\bm{y}|\bm{c})\neq p(\bm{y}|\bm{c},\bm{x})$, which is typically the case in real-world scenarios. In [16], information leakage is tackled from the perspective of out-of-distribution generalization. Concept Embedding Models (CEMs), introduced in [17], aim to alleviate the trade-off between accuracy and interpretability commonly observed in CBMs. However, we argue that this trade-off is inherently limited by the choice of concept set, and that CEMs may, in fact, be less interpretable than standard CBMs, as they yield more entropic representations—thereby amplifying information leakage. Other CEM-based approaches [18, 19, 20] are also prone to this issue. This critique is further elaborated in [21], which introduces an information-theoretic framework along with experimental benchmarks for evaluating interpretability in the representation space. Almost contemporaneously, [22] proposes a metric for interpretability grounded in the concept of usable mutual information [23].
Information Bottleneck in Representation Learning
The Information Bottleneck (IB) [12] is an information-theoretic framework that aims to balance the trade-off between preserving information about a factor variable and compressing the learned representation. A representation that retains all information about the factor is termed sufficient, while one that contains only information relevant to the factor—i.e., discarding all the irrelevant nuisances—is called minimal [24, 25, 26]. This terminology originates from the theory of minimal sufficient statistics [13, 14]. Direct computation of information-theoretic quantities is often intractable due to the complexity of the representation’s underlying distribution. To address this, variational approximations have been proposed to derive tractable evidence bounds that can be efficiently optimized [27, 28]. In this work, we leverage such approximations to optimize the IB objective with respect to the concepts.
Disentanglement
A closely related topic to CBMs that has been extensively studied is disentanglement. Despite its popularity, there remains no universally agreed-upon definition. However, existing definitions can be broadly categorized into three groups: (i) those in [8, 29, 30, 31, 32], which define a disentangled representation as one where changes in a specific part of the representation correspond to changes in a particular factor of variation, while remaining relatively invariant to other factors; (ii) those from works like [33, 34], which view a disentangled representation as one where modifying a single factor of variation leads to changes in only one component of the representation; and (iii) more recent perspectives [35, 36, 37], which consider a representation disentangled only if it satisfies both of the above properties. Conventional CBMs typically yield representations that align with the second definition. In contrast, MCBMs encourage representations to satisfy both criteria simultaneously by introducing an Information Bottleneck for each concept, thereby aligning with the third—and currently most widely accepted—definition of disentanglement. Our experiments demonstrate that MCBMs produce significantly more disentangled representations.
Binary - $\text{Bernoulli}\left(g^{y}_{\phi}(z)\right)$ $\text{Bernoulli}\left(g^{c}_{\phi}(z_{j})\right)$
Multiclass - $\text{Categoric}\left(g^{y}_{\phi}(z)\right)$ $\text{Categoric}\left(g^{c}_{\phi}(z_{j})\right)$
3 An Information-Theoretic Perspective on CBMs
@@ 3.2 Vanilla Models
In machine learning, the most commonly studied problem is that of predicting $\bm{y}$ from $\bm{x}$, which serves as a foundation for more specialized tasks. We refer to models trained to address this problem as Vanilla Models (VMs). These models typically operate by first extracting an intermediate representation $\bm{z}$ from the input $\bm{x}$ via an encoder $p_{\theta}(\bm{z}\mid\bm{x})$. Subsequently, a prediction $\hat{y}\in\mathcal{Y}$ is produced from $\bm{z}$ using a task head $q_{\phi}(\hat{\bm{y}}\mid\bm{z})$. Since $\bm{z}$ is intended to facilitate accurate prediction of $\bm{y}$, the mutual information between $\bm{z}$ and $\bm{y}$, denoted $I(Z;Y)$, should be maximized. In Appendix B.1, we formally show that:
@@
Figure 1(b) shows the graphical model of a Vanilla Model with a two-dimensional representation $\bm{z}$. Black edges represent the encoder $p_{\theta}(\bm{z}\mid\bm{x})$, while green edges indicate the task head $q_{\phi}(\hat{\bm{y}}\mid\bm{z})$.
The encoder is typically chosen to be deterministic, i.e., $p_{\theta}(\bm{z}\mid\bm{x})=\delta\left(\bm{z}-f_{\theta}(\bm{x})\right)$. However, for tractability reasons (see Section 4.2), we adopt a stochastic formulation where $p_{\theta}(\bm{z}\mid\bm{x})=\mathcal{N}\left(\bm{z};f_{\theta}(\bm{x}),\sigma_{x}^{2}I\right)$, as summarized in Table 1. The choice of task head $q_{\phi}(\hat{\bm{y}}\mid\bm{z})$ depends on the structure of the output space $\mathcal{Y}$, also detailed in Table 1. The objective in Equation (1) corresponds to minimizing the cross-entropy loss when $\bm{y}$ is binary or multiclass, and to minimizing the mean squared error between $\bm{y}$ and $g^{y}_{\phi}(z)$ when $\bm{y}$ is continuous.
@@ 3.3 Concept Bottleneck Models
As illustrated in Figure 1(c), CBMs extend Vanilla Models by incorporating a concept head $q(\hat{c}_{j}\mid z_{j})$, depicted with blue arrows. These models jointly optimize the objectives in Equations (1) and (2).
@@
As detailed in Table 1, the form of the concept head $q_{\phi}(\hat{c}_{j}\mid z_{j})$ depends on the nature of the concept space $\mathcal{C}$. The objective in Equation (2) corresponds to minimizing the cross-entropy loss when $c_{j}$ is binary or multiclass, and the mean squared error between $c_{j}$ and $g_{\phi}^{c}(z_{j})$ when $c_{j}$ is continuous.

### v2

@@ Abstract
Deep learning representations are often difficult to interpret, which can hinder their deployment in sensitive applications. Concept Bottleneck Models (CBMs) have emerged as a promising approach to mitigate this issue by learning representations that support target task performance while ensuring that each component predicts a concrete concept from a predefined set. In this work, we argue that CBMs do not impose a true bottleneck: the fact that a component can predict a concept does not guarantee that it encodes only information about that concept. This shortcoming raises concerns regarding interpretability and the validity of intervention procedures. To overcome this limitation, we propose Minimal Concept Bottleneck Models (MCBMs), which incorporate an Information Bottleneck (IB) objective to constrain each representation component to retain only the information relevant to its corresponding concept. This IB is implemented via a variational regularization term added to the training loss. As a result, MCBMs yield more interpretable representations, support principled concept-level interventions, and remain consistent with probability-theoretic foundations.
@@ 1 Introduction
Most machine learning models operate by learning data representations—compressed versions of the input that retain the essential information needed to solve a given task [1]. However, these representations often encode information in ways that are not easily interpretable by humans. This lack of interpretability becomes especially problematic in sensitive domains such as healthcare [2, 3, 4], finance [5, 6], and autonomous driving [7, 8]. To address this issue, Concept Bottleneck Models (CBMs) have been proposed, which enforce representations to be defined in terms of a set of human-understandable concepts [9].
Formally, given a task $\bm{y}$ and a set of concepts $\bm{c}=\{c_{j}\}_{j=1}^{m}$, Vanilla Models (VMs) are trained with a single objective: their representations $\bm{z}$ should capture the information necessary to predict $\bm{y}$ accurately. CBMs extend this framework by adding a second objective: each concept $c_{j}$ must be recoverable from a designated component $z_{j}\in\bm{z}$. By enforcing this additional constraint, CBMs are claimed to offer: (i) improved interpretability of the representation space, and (ii) the ability to intervene on specific concepts by modifying $z_{j}$ and propagating these changes to the predictions.
However, CBMs are prone to a phenomenon known as information leakage [10, 11], where the representation $\bm{z}$ encodes input information that cannot be attributed to the predefined concepts $\bm{c}$. We refer to this additional information as nuisances $\bm{n}$. Information leakage raises two main concerns: (i) it undermines interpretability, since $z_{j}$ cannot be fully explained by its corresponding concept $c_{j}$; and (ii) it compromises the validity of interventions—modifying $z_{j}$ may alter not only the associated concept $c_{j}$, but also other unintended information encoded in $z_{j}$.
@@
We argue that information leakage stems from a fundamental limitation in the current formulation of CBMs: the absence of an explicit Information Bottleneck (IB) [12] that actively constrains $z_{j}$ to exclude information unrelated to $c_{j}$. While the second objective in CBMs encourages each $z_{j}$ to retain $c_{j}$ in its entirety, it does not enforce that $z_{j}$ captures only information about $c_{j}$. In the worst-case scenario, $z_{j}$ could encode the entire input $\bm{x}$ and still satisfy this objective.
Table 1: Differences between VMs, CBMs and MCBMs.
VMs CBMs MCBMs
Does $z_{j}$ encode all $c_{j}$?
Does $z_{j}$ encode only $c_{j}$?
To address this issue, we propose Minimal Concept Bottleneck Models (MCBMs), which incorporate an IB into each $z_{j}$. This ensures that $z_{j}$ not only retains all the information about its associated concept $c_{j}$, but also excludes any information unrelated to $c_{j}$, as summarized in Table 1. The name reflects that $z_{j}$ is trained to be a minimal sufficient statistic of $c_{j}$ [13, 14], in contrast to traditional CBMs where $z_{j}$ is optimized to be merely a sufficient statistic of $c_{j}$. As illustrated in Figure 1, this design yields disentangled representations that directly address the two shortcomings of CBMs: (i) it improves interpretability, since $z_{j}$ can be fully explained by its corresponding concept $c_{j}$; and (ii) it enables valid interventions—modifying $z_{j}$ affects only the associated concept $c_{j}$.
In Section 2, we connect the data generative process to MCBMs through information-theoretic quantities, showing that the IB can be implemented via a variational loss. In Section 3, we review alternative approaches to address information leakage. In Section 4, we present experiments demonstrating that MCBMs enforce a true bottleneck, thereby enhancing interpretability and intervenability compared to existing alternatives. Finally, in Section 5, we show that the assumptions made in CBMs to enable interventions are theoretically flawed.
Figure 1: In Vanilla Models, concepts and nuisances may be arbitrarily entangled in the representation, and a variable $z_{j}$ may capture only part of a concept (depicted as paler colors). In CBMs, each $z_{j}$ encodes all information about its corresponding concept $c_{j}$, but may also capture some information about nuisances (e.g., $z_{1}$) or other concepts (e.g., $z_{2}$). In contrast, MCBMs enforce that each representation variable $z_{j}$ encodes all—and only—the information about its corresponding concept.
2 From Data Generative Process to MCBMs
@@ 2.2 Vanilla Models
In machine learning, the most commonly studied problem is that of predicting $\bm{y}$ from $\bm{x}$, which serves as a foundation for more specialized tasks. We refer to models trained to address this problem as Vanilla Models (VMs). These models typically operate by first extracting an intermediate representation $\bm{z}\in\mathcal{Z}$ from the input $\bm{x}$ via an encoder $p_{\theta}(\bm{z}\mid\bm{x})$. Subsequently, a prediction $\hat{y}\in\mathcal{Y}$ is produced from $\bm{z}$ using a task head $q_{\phi}(\hat{\bm{y}}\mid\bm{z})$. Since $\bm{z}$ is intended to facilitate accurate prediction of $\bm{y}$, the mutual information between $\bm{z}$ and $\bm{y}$, denoted $I(Z;Y)$, should be maximized. In Appendix B.1, we formally show that:
@@
Figure 2(b) shows the graphical model of a Vanilla Model with a two-dimensional representation $\bm{z}$. Black edges represent the encoder $p_{\theta}(\bm{z}\mid\bm{x})$, while green edges indicate the task head $q_{\phi}(\hat{\bm{y}}\mid\bm{z})$. The encoder is typically chosen to be deterministic, i.e., $p_{\theta}(\bm{z}\mid\bm{x})=\delta\left(\bm{z}-f_{\theta}(\bm{x})\right)$. However, for tractability reasons (see Section 2.4), we adopt a stochastic formulation where $p_{\theta}(\bm{z}\mid\bm{x})=\mathcal{N}\left(\bm{z};f_{\theta}(\bm{x}),\sigma_{x}^{2}I\right)$, as summarized in Table 2. The choice of task head $q_{\phi}(\hat{\bm{y}}\mid\bm{z})$ depends on the structure of the output space $\mathcal{Y}$, also detailed in Table 2. The objective in Equation 1 corresponds to minimizing the cross-entropy loss when $\bm{y}$ is binary or multiclass, and to minimizing the mean squared error between $\bm{y}$ and $g^{y}_{\phi}(z)$ when $\bm{y}$ is continuous.
@@ 2.3 Concept Bottleneck Models
As illustrated in Figure 2(c), CBMs extend Vanilla Models by incorporating a concept head $q(\hat{c}_{j}\mid z_{j})$, depicted with blue arrows. These models jointly optimize the objectives in Equations 1 and 2. As detailed in Table 2, the form of the concept head $q_{\phi}(\hat{c}_{j}\mid z_{j})$ depends on the nature of the concept space $\mathcal{C}$. The objective in Equation 2 corresponds to minimizing the cross-entropy loss when $c_{j}$ is binary or multiclass, and the mean squared error between $c_{j}$ and $g_{\phi}^{c}(z_{j})$ when $c_{j}$ is continuous.
How are Interventions Performed in CBMs?
As explained in Section 1, a key advantage often attributed to CBMs is their ability to support concept-level interventions. Suppose we aim to estimate $p(\hat{\bm{y}}\mid c_{j}=\alpha,\bm{x})$. In CBMs, this intervention is performed through the latent representation $z_{j}$:
[…]
However, the conditional distribution $p(z_{j}\mid c_{j})$ is not defined—there is no directed path from $c_{j}$ to $z_{j}$ in Figure 2(c). Intuitively, because $z_{j}$ may encode information about $\bm{x}$ beyond $c_{j}$, it cannot be fully determined by $c_{j}$ alone. This raises a key question: how can interventions be performed in CBMs if $p(z_{j}\mid c_{j})$ is unknown? To make interventions feasible, CBMs typically impose two constraints:
  1. (i)
    Concepts $c_{j}\in\bm{c}$, are assumed to be binary. If a concept is originally multiclass with $k$ categories, it is converted into $k$ binary concepts (One-vs-Rest [15]).
  2. (ii)
Since $\sigma$ is invertible, this setup permits defining $p(z_{j}\mid c_{j})\approx\sigma^{-1}(c_{j})$. However, this is ill-defined at the binary extremes, as $\sigma^{-1}(1)=-\sigma^{-1}(0)=\infty$. To address this, in practice, $p(z_{j}\mid c_{j}=0)$ and $p(z_{j}\mid c_{j}=1)$ are set as the 5th and 95th percentiles of the empirical distribution of $z_{j}$, respectively [9]. This workaround, however, introduces two crucial issues discussed in Section 5.
Binary - $\text{Bernoulli}\left(g^{y}_{\phi}(z)\right)$ $\text{Bernoulli}\left(g^{c}_{\phi}(z_{j})\right)$ -
Multiclass - $\text{Categoric}\left(g^{y}_{\phi}(z)\right)$ $\text{Categoric}\left(g^{c}_{\phi}(z_{j})\right)$ -
2.4 Minimal Concept Bottleneck Models
As discussed in Section 1, CBMs lack an explicit mechanism for enforcing a bottleneck, which often undermines their intended advantages. To address this limitation, we introduce Minimal Concept Bottleneck Models (MCBMs), which explicitly impose an Information Bottleneck. This ensures that each $z_{j}$ retains all information about its associated concept $c_{j}$, while excluding information unrelated to $c_{j}$. In other words, $z_{j}$ becomes a minimal sufficient representation of $c_{j}$. To achieve this, we introduce a representation head $q_{\phi}(\hat{z}_{j}\mid c_{j})$ that predicts $z_{j}$ from $c_{j}$. This encourages $z_{j}$ to discard information unrelated to $c_{j}$, as doing so improves the predictive accuracy of $q_{\phi}(\hat{z}_{j}\mid c_{j})$. Formally, this objective corresponds to minimizing the conditional mutual information $I(Z_{j};X\mid C_{j})$. We note that if $z_{j}$ is a representation of $\bm{x}$, then the following propositions are equivalent: (i) $I(Z_{j};X|C_{j})=0$, (ii) the Markov Chain $X\leftrightarrow C_{j}\leftrightarrow Z_{j}$ is satisfied and (iii) $p(z_{j}|c_{j})=p(z_{j}|x)$. To minimize $I(Z_{j};X\mid C_{j})$, we leverage the following identity, which is proven in Appendix B.3:
Figure 2(d) illustrates how MCBMs extend CBMs by introducing the representation head $q(\hat{z}_{j}\mid c_{j})$, depicted with red arrows. These models are trained to jointly optimize the objectives given in Equations 1, 2, and 4. Although the KL Divergence in Equation 4 does not admit a closed-form solution in general, we show in Appendix B.4 that, under the distributional assumptions listed in Table 2, it reduces to the mean squared error between $f_{\theta}(x)$ and $g^{z}_{\phi}(c_{j})$.
How are Interventions Performed in MCBMs?
In contrast to CBMs, MCBMs explicitly constrain $z_{j}$ to contain only information about $c_{j}$. As a result, modifying $z_{j}$ corresponds to intervening solely on $c_{j}$. This is reflected in Figure 2(d), where MCBMs introduce a directed path from $c_{j}$ to $z_{j}$ through the intermediate variable $\hat{z}_{j}$. This structure permits the computation of:
$p(z_{j}|c_{j})=\int p(z_{j}|\hat{z}_{j})q_{\phi}(\hat{z}_{j}|c_{j})\,d\hat{z}_{j}$ (5)
which simplifies to $p(z_{j}\mid c_{j})=q_{\phi}(z_{j}\mid c_{j})$ when $p(z_{j}\mid\hat{z}_{j})=\delta(z_{j}-\hat{z}_{j})$, i.e., once the objective in Equation 4 is optimized and $z_{j}$ encodes exclusively $c_{j}$. As shown in Table 2, we define the encoder distribution as $q_{\phi}(z_{j}\mid c_{j})=\mathcal{N}(g^{z}_{\phi}(c_{j}),\sigma_{\hat{z}}^{2}I)$, where mean function $g^{z}_{\phi}(c_{j})$ is chosen according to the following rules:
  1. (i)
  2. (ii)
    For categorical concepts, $g^{z}_{\phi}(c_{j})=\lambda\cdot\mathrm{one\_hot}(c_{j})$. This mirrors the Prototypical Learning [16] approach where class-dependent prototypes $\{g^{z}_{\phi}(c_{j})\}_{j=1}^{m}$ are fixed.
  3. (iii)
    For continuous concepts, $g^{z}_{\phi}(c_{j})=\lambda\cdot c_{j}$.
Here, $\lambda$ is a scaling constant that controls the norm of the latent representation, fixed to $\lambda=3$ in all experiments. Regarding the variance term, we set: (i) $\sigma_{x}=0$in the case of CBMs to obtain a deterministic encoder in line with their original formulation, and (ii) $\sigma_{x}=\sigma_{\hat{z}}=1$for MCBMs.
Practical Considerations for Optimizing MCBMs
To optimize MCBMs, we combine the three previously introduced objectives, incorporating two key considerations. First, we approximate the expectations over $p(\bm{x},\bm{y})$ and $p(\bm{x},c_{j})$ using the empirical data distribution, replacing integrals with summations over the dataset. Second, to enable gradient-based optimization through the stochastic encoder, we apply the reparameterization trick [17]: $\mathbb{E}_{p_{\theta}(\bm{z}|\bm{x})}\left[\log q_{\phi}(\bm{y}|\bm{z})\right]\approx\sum_{i}\log q_{\phi}\left(\bm{y}|f^{\prime}_{\theta}\left(x,\epsilon^{(i)}\right)\right)$ and $E_{p_{\theta}(z_{j}|\bm{x})}\left[\log q_{\phi}(\hat{c}_{j}|z_{j})\right]\approx\sum_{i}\log q_{\phi}\left(\hat{c}_{j}\big|f^{\prime}_{\theta,j}\left(\bm{x},\epsilon^{(i)}\right)\right)$, where $f^{\prime}_{\theta}(\bm{x},\epsilon)=f_{\theta}(\bm{x})+\sigma^{2}_{x}I\epsilon$ (due to the choice of $p_{\theta}(\bm{z}|\bm{x})$ in Table 2), $\epsilon\sim\mathcal{N}\left(0,I\right)$ and $f^{\prime}_{\theta,j}\left(x,\epsilon\right)$ corresponds to the element $j$ of $f^{\prime}_{\theta}\left(\bm{x},\epsilon\right)$. Combining the considerations above yields the final objective for training MCBMs, as shown in Equation 6, where $\beta$ and $\gamma$ are hyperparameters. The first term corresponds to the objective used in Vanilla Models, the second term is introduced in CBMs, and the third is specific to MCBMs. Detailed training algorithms for the various cases listed in Table 2 are provided in Appendix C.